OpenStudy (anonymous):

solve the equation for x. identify any extraneous solutions. √x = -2

OpenStudy (solomonzelman):

square both sides

OpenStudy (anonymous):

@SolomonZelman both sides squared would be x=-4

OpenStudy (anonymous):

@SolomonZelman ??

OpenStudy (dan815):

(-2) ^2 = -2 * -2

OpenStudy (dan815):

are you sure its -4?

OpenStudy (dumbcow):

This has no solution, the sqrt of a number must be greater than 0

OpenStudy (anonymous):

@dan815 um i just typed in -2^2 on my calculator and it gave me -4.??

OpenStudy (dumbcow):

(-2)^2 ;)

OpenStudy (dan815):

type (-2)^2

OpenStudy (dan815):

-2^2 = -1*2^2 which is not the same as-2*-2 = -1*2*-1*2

OpenStudy (anonymous):

@dumbcow @dan815 okay so 4 would be a solution? because i typed in (-2)^2 on my calculator and it gave me 4.

OpenStudy (solomonzelman):

square (cube or any other) root of a number doesn't have to be greater than 0.

OpenStudy (dan815):

yes but i hope this makes sense just remembeer for now that -1 * -1 = 1

OpenStudy (amistre64):

by default, there is no real value that gets us to sqrt(x)= -2 since the range of sqrt(x) 0 and above: